What is the value of 1+ sin theta -cos theta /1+cos theta +sin theta

Grade:8

1 Answers

Arun
6 years ago
Dear Student
 
[(1 + sin theta) - cos theta]² / [(1+ sin theta)² - cos² theta] 
= 1 + sin² theta + 2 sin theta + cos² theta - 2 sin theta* cos tgeta - 2 cos theta / 1 + sin² theta + 2 sin theta - cos² tgeta
= 2 ( 1+ sin theta) -2 cos theta(1 + sin theta) / 2 sin theta (1 + sin theta)
= 1 - cos theta/ sin theta
 
Regards
Arun (askIITians forum expert)